# Automatic Scoring of Students’ Science Writing Using Hybrid Neural Network

### 02 Dec 2023 v1

@@ 7 Discussion and Conclusions
In conclusion, the suggested multi-perceptive hybrid neural network approach offers a potentially effective means of addressing the difficulties associated with evaluating complex constructs, such as the knowledge-in-use skills required for scientific practice as described in the Framework for K–12 Science Education. The results of the current study, which assessed the suggested approach’s effectiveness and put it up against other conventional ML models, showed that the multi-perceptive approach was more accurate in determining the different perceptual expectations of NGSS tasks. Furthermore, by highlighting the possibilities of using label semantics and fine-grained text data to create a categorized scoring scheme that narrates a student’s cognitive process for a particular assignment, this study contributes to the field of automated assessment in education. Future studies will examine how well the suggested strategy may be applied in other educational settings and further develop the methodology to improve its real-world applications.

### 26 Dec 2023 v2

@@ 7 Discussion and Conclusions
In conclusion, the suggested multi-perspective hybrid neural network approach offers a potentially effective means of addressing the difficulties associated with evaluating complex constructs, such as the knowledge-in-use skills required for scientific practice as described in the Framework for K–12 Science Education. The results of the current study, which assessed the suggested approach’s effectiveness and put it up against other conventional ML models, showed that the multi-perspective approach was more accurate in determining the different perceptual expectations of NGSS tasks. Furthermore, by highlighting the possibilities of using label semantics and fine-grained text data to create a categorized scoring scheme that narrates a student’s cognitive process for a particular assignment, this study contributes to the field of automated assessment in education. Future studies will examine how well the suggested strategy may be applied in other educational settings and further develop the methodology to improve its real-world applications.